OpenStudy (anonymous):

Will give medal to correct answer! :3 Solve, using linear combination. 5x + y = 2 4x + y = 4 A.) (–3, 17) B.) (–2, 12) C.) (0, 2) D.) (1, –8)

OpenStudy (anonymous):

(5x+y)-(4x+y)=x=2-4=-2; so answer must be B.

OpenStudy (anonymous):

Thank you!